Calculate Kb for the conjugate base of carbonic acid, H₂CO₃(aq), Ka = 4.4x10⁻⁷
1
Expert's answer
2020-06-15T14:41:27-0400

The value of base ionisation constant "Kb" is calculated using the following formula,

"Kb=\dfrac{Kw}{Ka} where,"

"Ka" is the acid dissociation constant

"Kw" is water dissociation constant

We are given,

"Ka=" "4.4*10^{-7}"

"Kw=1.0*10^{-14}"

Substituting

"Kb=\dfrac{1.0*10^{-14}}{4.4*10^{-7}}"


"=2.3*10^{-8}"

"Kb" for conjugate base of carbonic acid"=2.3*10^{-8}"